Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
takahiro.yamamoto - 7:51 Monday 17 February 2020 (12994) Print this report
Comment to CARM loop cleaned up (12961)

Estimated sensitivity had been worse as a factor of 1.5 because the exchange of the coil driver chassis had been missed in the calibration.
(Super high power coil driver had been used for ETMY TM.)

Calibration and online h(t) reconstruction was fixed as I reported on klog12992.
Unfortunately, IFO lock is difficult today due to the microseismic disturbance, but correct BNS range can be estimated as roughly 600kpc.

Search Help